Given the following end of year prices and states of the economy. Combine the following assets to form a portfolio with zero risk. What is the weight of AAA in your portfolio? What is the expected return of the portfolio? Enter your answer as a percent. Round your answer to two decimals.

Probability

.25

.75

Price today

Expansion

Recession

AAA

100

130

110

BBB

50

55

65
Answer #1

ANSWER

Working note 1

Calculation of price of AAA at end of year

State of Economy Probability(a) Price(b)

Expected

Price(c=a*b)

Expansion 0.25 130 32.50
Recession 0.75 110 82.50
TOTAL 115

Working note 2

Calculation of price of BBB at end of year

State of Economy Probability(a) Price(b)

Expected

Price(c=a*b)

Expansion 0.25 55 13.75
Recession 0.75 65 48.75
TOTAL 62.50

Working note 3

Calculation of return on AAA

Return = (Price at end- Price today) / Price today

= (115-100) / 100

= [15/100]*100

Return = 15%

Working note 4

Calculation of return on BBB

Return = (Price at end - Price today) / Price today

= (62.50 - 50) / 50

= [12.50 / 50]*100

Return = 25%

Calculation of weight of AAA and return of portfolio

Security

Price at end

of year (a)

Weight

(b= a/177.50)

Return

(c)

Weight * Return

(d= b*c)

AAA 115 0.65 15% 9.75%
BBB 62.50 0.35 25% 8.75%
TOTAL 177.50 1.00 18.50%

Hence,

1) Weight of AAA in portfoli = 65% (approx.)

2) Return on portfolio = 18.50%